OpenStudy (anonymous):

Simplify. Identify any x-values for which the expression is undefined. \[\frac{ -x ^{2}+3x }{ 2x ^{2}-7x+3 }\]

OpenStudy (anonymous):

for the undefined : 2x^2 -7x + 3 =0 2x^2 -6x-x + 3 =0 2x(x-3)-1(x-3) = 0 (2x-1)(x-3)=0 so x=3 and x = 1/2

OpenStudy (anonymous):

-x/(2x-1)

OpenStudy (anonymous):

the numerator is -x^2 + 3x = x(3-x) so you have : x(3-x) / (2x-1)(x-3) the division of (3-x) and (x-3) gives -1 so you left with : -x / (2x-1) or better : x/ (1-2x)

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

smalls could you start trying to do these problems then you post what you have and we will check it and see where you get difficulties, this way you can get better grades on the tests

OpenStudy (anonymous):

surdawi - you better not give the final answer .. its better to write how you got there

OpenStudy (anonymous):

i only have like two more questions i just have class in an hour and a half and kind of get how to do it but just need my work done but i have a meeting with her after to tutor me